Find the vertex, focus, and directrix of the parabola given by the equation. Sketch the graph.
(2x − 8)2 = 8y − 56

Answer by ewatrrr(24785)

 
Hi,
(2x - 8)^2 = 8y- 56
2x^2 - 16x + 64 = 8y - 56
2(x-4)^2 -32 +64 = 8y - 56
2(x-4)^2 = 8y - 88
2(x-4)^2 = 8(y-11)
(x-4)^2 = 4(y-11) 4p = 4, p = 1, V(4,11), F(4,12) directrix: y = 10
y = (1/4)(x-4)^2 + 11 line of Symmetry is x = 4